OpenStudy (waheguru):

Dont understand this question - Identify equal rational numbers in the list that follows 2/3 -3/2 -2/3 2 / 3 (a negative sign beside the line separating the 2 numbers) 3/2 (a negative sign beside the line separating the 2 numbers) 2/-3 3/-2 and 3/2

OpenStudy (waheguru):

here is the question

OpenStudy (anonymous):

why do people repost questions that someone's helping them with?

OpenStudy (waheguru):

i dont get it

OpenStudy (waheguru):

then how is 2/3 equal to the rest of the numbrs

OpenStudy (anonymous):

2/3 doesn't...the question doesn't say that they *all* equal something else. just to find the ones that *are* equal.

OpenStudy (waheguru):

whait ill show u the answer

OpenStudy (anonymous):

you have 8 fractions there...thre of them we've established are equal to each other, right?

OpenStudy (waheguru):

ok so can i write them in any order

OpenStudy (anonymous):

right...that's what i said. we've established the first set of three. The second set of three (separated by the semi-colon) are the others that are the same. 3/2 and 2/3 from your origianl list aren't equal to any others. order doesn't matter a=b=c is the same as a=c=b or c=a=b

OpenStudy (waheguru):

ok so can i write them in any order

OpenStudy (waheguru):

oh i got it

OpenStudy (anonymous):

the idea here is that you understand that \[\frac{-a}{b}=\frac{a}{-b}=-\frac{a}{b}\]

OpenStudy (anonymous):

that's all they are getting you to recognize :)

OpenStudy (waheguru):

thankyou soo much bro

OpenStudy (anonymous):

np :)

OpenStudy (waheguru):

=)